Assume that random guesses are made for 8 multiple choice questions on a SAT test. Each question has 5 possible answers,

What is the probability of getting a question correct?
Answer #1

Assume that random guesses are made for 8 multiple choice questions on a SAT test. Each question has 5 possible answers

So here for each questions random guesses are made.

only one of 5 possible answer is correct.

Therefore,

Probability of getting a question correct is = (no. of correct answer)/(total possible answer) = 1/5=0.2

Thus

probability of getting a question correct is 0.2